Product of open loop gain & its bandwidth =10^6 Hz
Calculate close loop bandwidth of amplifier .given R(I)=10k ohm
&R(f)=150k ohm
1
Expert's answer
2019-07-09T13:14:17-0400

For an amplifier the product of gain and bandwidth is constant.

For a non-inverting amplifier, the closed loop gain ("A_{CL}") is given by

"A_{CL}=1+\frac{R_f}{R_i}"

"A_{CL}=1+\frac{150}{10}"

"A_{CL}=16"

If closed loop bandwidth is given by "B_{CL}" then

                                             "A_{CL}B_{CL}="  Product of open loop gain & its bandwidth

"A_{CL}B_{CL}=10^6"

"16B_{CL}=10^6"

"B_{CL}=6.25\times 10^4 \ Hz"

Answer: The closed loop bandwidth of the amplifier is  "6.25\times 10^4\ Hz"